# AtomWorld: A Benchmark for Evaluating Spatial Reasoning in Large Language Models on Crystalline Materials

### v2

@@ 1 Introduction
In this work we introduce the AtomWorld Benchmark to evaluate reasoning LLMs on CIF motor skills. To the best of our knowledge, we are the first benchmark to examine this fundamental skill of crystallography in LLMs. Moreover, AtomWorld at its core is a scalable data generator, which can be used to support LLM training. While these tasks can be manually solved via software or packages such as Ovito and Atomic Simulation Environment (ASE) (Stukowski, 2010; Hjorth Larsen et al., 2017), installing this capability in LLMs is necessary for the more valuable downstream cognitive tasks, e.g. building an agentic workflow for material discovery.
@@
We evaluated our benchmark across several frontier text-based models, which through algorithmic approaches were generally able to succeed at simpler tasks such as adding or moving atoms, but struggled with more complex tasks such as rotating around an atom. Notably, some actions intuitive to humans had high error rates, while others considered more tedious were solved unexpectedly well. The AtomWorld benchmark is also complemented through a series of tests (PointWorld, CIF-Gen, CIF-Repair, Chemical Competence Score, Struct-Prop) to isolate different aspects of LLM weaknesses and explore the relation that motor skills tests have with perceptual and cognitive skills. Traditionally, LLMs have struggled with the spacial reasoning and long syntax following skills required by our benchmark - but this may soon change with rapid advancements in tool-augmented design (Hu et al., 2025), diffusion LLMs (Nie et al., 2025; Song et al., 2025), and as language-aligned video generation (Zheng et al., 2024; DeepMind, 2025) and robotics (Assran et al., 2025) models become increasingly capable. We hope that our AtomWorld playground can play a foundational role in both testing and developing the understanding of 3D CIF environments in tomorrow’s LLMs.
@@ 2.1 AtomWorld Generator
At its core, AtomWorld is designed as a data generator. This data follows a three-part structure: two CIF files of “before” and “after” states, and an action prompt describing the change - with the goal of the LLM to yield the “after” state, given the “before” state and action. A flowchart describing the workflow from data generator to benchmark is presented in Figure 1.
Figure 1: AtomWorld benchmark flowchart. The AtomWorld generator follows a structured data flow: the random sampler selects a structure from a predefined structure pool (in this work, a subset of CIF files from the Materials Project database (Jain et al., 2013)); the random initializer parametrizes the chosen action template by assigning atom indices and/or positions; the structure operator applies the instantiated action to the original structure to obtain the target structure; and the prompter generates a natural language description aligned with the action. The resulting (input structure, action prompt) pairs are then fed into the LLM agent system, whose generated structure is compared against the target structure using the StructureMatcher from pymatgen (Ong et al., 2013) to compute the desired evaluation metric.
@@
swap Swap atoms at indices {index1} and {index2} in the cif file. The indices of atoms are started from 0.
@@ 2.2 AtomWorld for LLM Training
2.3 Complementary Benchmarks
Beyond AtomWorld, we design a set of benchmarks to probe different aspects of LLMs’ structural reasoning with CIF data. Together, these tasks complement AtomWorld to form a comprehensive suite that spans format literacy, spatial reasoning, and property-oriented understanding.
@@ 2.3 Complementary Benchmarks
      CIF-Gen: Evaluates whether the model can explicitly produce syntactically valid CIFs for simple prototype crystals (e.g., sc, fcc, bcc, perovskite), thereby examining its familiarity with CIF conventions and basic materials knowledge, as opposed to open-ended CIF generation.
@@
      Chemical Competence Score (CCS): This test assesses a model’s latent chemical knowledge by evaluating its precision in distinguishing chemically accurate from inaccurate descriptions of crystal structures. While this test is a “perceptual skills” test, we use this to measure the effect that chemistry pretraining has on LLM performance in “motor skills” tasks. Following the methodology of Bran et al. (2025), the dataset was constructed by sampling 600 unique crystal structures from the Materials Project, with corresponding descriptions generated using Robocrystallographer (Ganose & Jain, 2019). An inaccurate dataset was then created by replacing one sentence in each original description with a sentence describing a different crystal. Because the CCS is computed from the token log-likelihoods at the model’s final layer, access to these probabilities is required; this score can be calculated only for locally-run models.
@@ 3.1 Models and Parameter Ranges
Our selection of LLMs covers frontier closed models and strong open-source baselines. We chose the Qwen-3 series to test for parameter scaling effects. We also considered science-specialised LLMs; e.g. NatureLM (Xia et al., 2025), and MatterChat (Tang et al., 2025). However, these were excluded due to either unable to produce outputs in the required CIF format, or not currently accessible via public APIs or code implementations. For tool-augmented LLMs, we designed a preliminary framework that enables interaction with tools such as Pymatgen. Although this workflow still requires refinement and future iterations may yield stronger results, the current findings already illustrate meaningful trends.
@@ 3.3 Datasets
    AtomWorld. Contains around 1000 samples for each of the 10 action types. Due to testing cost considerations, we evaluated a subset of samples for each model: 250 samples for actions add, move, move_towards, insert_between, and rotate_around; and 50 samples for actions remove, change, swap, delete_below, and super_cell. For the super_cell action, the output structure was specified to range from 2 to 8x the original cell size. An example of an insert_between test is illustrated in Appendix A.3.
@@ 4 Results
4.1 AtomWorld
The main results of the AtomWorld benchmark, alongside complementary tests are presented in Figure 2.a. We see some separation of the AtomWorld actions into easy (change, remove, add), moderate (move, move towards, insert between) and hard difficulty (swap, delete below, rotate around) levels based on their success rates. The moderate and hard difficulty tasks constitute greater requirements of multi-step or spatial reasoning. We also notice the mean max_dist metric increase for the more difficult tasks (minus tasks not requiring structural perturbations). One interesting finding is that the swap action was intended to be easy, since it only requires an exchange of two element types, or their coordinates. But all of the models, especially the larger ones, failed surprisingly. Also notably, the super_cell task cannot be well categorised into these difficulty tiers as the success rates range from Llama3-70B’s 4% to 98% from GPT-o3 - it’s both easy (just large-scale repetition) and difficult (requires long-context output) at the same time. The parameter scaling results in Figure 2.c and d illustrate that larger models generally achieve higher success rates and smaller displacements. However, with improvements with scale being marginal with more difficult tasks, and noting that Qwen3-32B outperformes Llama3-70B across most tasks, it suggests that architectural design and training strategies play an equally important role as parameter size.
Our evaluation of our tool-augmented LLM framework on AtomWorld tasks found noticeable gains in model performance. However, the gains are somewhat limited, particularly for more complex actions. Detailed results and comprehensive analysis can be found in Appendix B.1.
Figure 2: a. Success rate metric across AtomWorld, CIF-Repair, CIF-Gen and StructProp datasets. b. Mean max_dist metric across AtomWorld and CIF-Gen datasets. c, d. Parameter scaling results on Qwen3 series.
@@ 4.2 PointWorld and CIF Literacy Tests
Figure 3: The number of correctly generated CIFs for each structure type in the CIF-Gen task. The squares marked in red indicate cases where the single correct generation is the standard prototype. The right side shows the specific 3D crystal structures for each type, where the chemical compositions in red represent the standard prototypes.
@@ 5 Discussion
6 Related Work
LLMs for crystallography.
LLMs have been primarily explored for their capabilities in CIF generation and QA. LLMs have been demonstrated to hold an innate ability to generate crystal structures when pretrained on millions of CIF files (Antunes et al., 2024). This process may be further reinforced through evolutionary search frameworks (Gan et al., 2025). However, as LLMs are pattern predictors, the search space is fundamentally limited by the scope of the pretraining data. LLMs can also be instruction fine-tuned to predict crystal properties or provide general QA responses from CIF, e.g. AlchemBERT, NatureLM, Darwin 1.5, etc (Liu et al., 2025; Xia et al., 2025; Xie et al., 2025; Van Herck et al., 2025; Nate Gruver & Ulissi, 2024). Crystallography QA is well benchmarked, with the most comprehensive being LLM4Mat-Bench (Niyongabo Rubungo et al., 2025), consisting of approximately 2 million composition-structure-description pairs. Tool-augmented LLMs such as OSDA Agent (Hu et al., 2025) improve structure generation through coupling computational chemistry tools to LLMs. These tool-augmented design frameworks are able to address the lack of in-depth chemistry knowledge of LLMs without expensive (and not always effective) fine-tuning. LLMs may be able to reliably handle geometric CIF modification through tool-augmentation.
Multimodal reasoning.
@@ Appendix A AtomWorld Setup Details
A.1 Supported action prompts
@@
A.2 Full Prompt Templates
@@ 2. Lattice Parameter Adjustment;
A.3 Illustrative example of the framework
@@
A.4 Logic on generating CIF-Repair task
@@
A.5 DFT computation details
All density functional theory (DFT) calculations, including band gap and bulk modulus evaluations, were performed using the Vienna Ab initio Simulation Package (VASP) with the projector-augmented wave (PAW) method (Kresse & Hafner, 1993; Kresse & Furthmüller, 1996a; b; Kresse & Joubert, 1999) and the PBEsol exchange–correlation functional(Perdew et al., 2008). High-throughput workflows for both properties were automated using the atomate2 package (Ganose et al., 2025). Unless otherwise specified, calculation parameters followed the default settings in atomate2. Example calculation scripts are provided in the github repository.
@@ A.5 DFT computation details
Appendix B Full Evaluation Results
B.1 Evaluations of tool augmented LLM for AtomWorld

### v3

@@ 1 Introduction
LLMs for material discovery would primarily benefit researchers at the cognitive stage. Several works such as CrystaLLM, MatterGPT, and AtomGPT (Antunes et al., 2024; Chen et al., 2024; Choudhary, 2024) aim to leverage the generative capability of LLMs to generate useful and novel CIFs. These works however, are fundamentally limited by the scope of pretraining data - generations are plausible mostly only for data-rich domains of ideal, bulk crystals. The hypothesis which motivates our work is that LLMs could generate useful CIFs past data-rich domains if the more tractable problem of CIF modification is solved as opposed to ab initio generation. Learning CIF modification is learning how rather than what to generate, requiring both motor and perceptual skills as a prerequisite. In current literature, perceptual skills have been tested through question-answer (QA) style benchmarks e.g. LLM4Mat-Bench (Niyongabo Rubungo et al., 2025), but less attention has been given to testing motor skills. To address this gap, our research question asks: how can we measure and improve LLM “crystallographic motor skills”, i.e. ability to manipulate atoms in crystal structures?
@@
In this work we introduce AtomMotor-1K, a compact test set of 1500 questions to benchmark reasoning LLMs on CIF motor skills. The test is made of 10 CIF action types, broken down from the real-world structural modifications which researchers may perform on CIFs. To the best of our knowledge, this is the first benchmark which examines this fundamental skill of crystallography in LLMs. The dataset of AtomMotor-1K was generated from our AtomWorld data generator, which can generate unlimited samples for each action, and also be used to support LLM training. We used AtomMotor-1K to evaluate several frontier text-based models, which through algorithmic approaches were generally able to succeed at simpler tasks such as adding or moving atoms, but struggled with more complex tasks such as rotating around an atom. Notably, some actions intuitive to humans had high error rates, while others considered more tedious were solved unexpectedly well. Furthermore, for the analysis of AtomMotor-1K we designed a series of tests (PointWorld, CIF-Gen, CIF-Repair, Chemical Competence Score, Struct-Prop) to isolate different aspects of LLM weaknesses and explore the relation that motor skills tests have with perceptual and cognitive skills.
We expect AtomWorld and AtomMotor-1K to gain increasing relevance as LLMs improve. LLMs have traditionally struggled with the spacial reasoning and long syntax following skills required by our benchmark - but this may soon change with rapid advancements in tool-augmented design (Hu et al., 2025), diffusion LLMs (Nie et al., 2025; Song et al., 2025), and as language-aligned video generation (Zheng et al., 2024; DeepMind, 2025) and robotics (Assran et al., 2025) models become increasingly capable. LLMs capability to manipulate crystal geometries is an important yet underexplored topic, and we believe the AtomWorld playground can play a foundational role in both testing and developing this in tomorrow’s LLMs.
2 Related Work
LLMs for crystallography.
LLMs have been primarily explored for their capabilities in CIF generation and QA. LLMs have been demonstrated to hold an innate ability to generate crystal structures when pretrained on millions of CIF files (Antunes et al., 2024). This process may be further reinforced through evolutionary search frameworks (Gan et al., 2025). However, as LLMs are pattern predictors, the search space is fundamentally limited by the scope of the pretraining data. LLMs can also be instruction fine-tuned to predict crystal properties or provide general QA responses from CIF, e.g. AlchemBERT, NatureLM, Darwin 1.5, etc (Liu et al., 2025; Xia et al., 2025; Xie et al., 2025; Van Herck et al., 2025; Nate Gruver and Ulissi, 2024). In data modelling, MatText (Alampara et al., 2025) investigates if such QA can be improved through different textual representations. Crystallography QA is well benchmarked, with the most comprehensive being LLM4Mat-Bench (Niyongabo Rubungo et al., 2025), consisting of approximately 2 million composition-structure-description pairs. Crystallography benchmarks also cover multimodal LLMs, e.g. work by Polat et al. (2025) investigates the generation of structural annotations to crystallographic images.
Tool-augmented LLMs such as OSDA Agent (Hu et al., 2025) improve structure generation through coupling computational chemistry tools to LLMs. These tool-augmented design frameworks are able to address the lack of in-depth chemistry knowledge of LLMs without expensive (and not always effective) fine-tuning. LLMs may be able to reliably handle geometric CIF modification through tool-augmentation.
Multimodal reasoning.
@@ 3.1 AtomWorld Generator
The dataset of AtomMotor-1K was generated from the scalable AtomWorld data generator. The data follows a three-part structure: two CIF files of “before” and “after” states, and an action prompt describing the change - with the goal of the LLM to yield the “after” state, given the “before” state and action. A flowchart describing the workflow from data generator to benchmark is presented in Figure 1.
The CIF format contains many optional and extensible fields. Using arbitrarily mixed CIF variants would introduce additional sources of uncertainty that are not directly related to the manipulation abilities we intend to evaluate. For this reason, we adopt the default CIF representations generated from pymatgen (Ong et al., 2013) and the Materials Project (MP) (Jain et al., 2013) as a standardized input format for all tasks. We leave studying how LLM performance varies across different CIF styles, as well as across other structure formats such as POSCAR and XYZ for future work.
Figure 1: AtomWorld benchmark flowchart. The AtomWorld generator follows a structured data flow: the random sampler selects a structure from a predefined structure pool; the random initializer parametrizes the chosen action template by assigning atom indices and/or positions; the structure operator applies the instantiated action to the original structure to obtain the target structure; and the prompter generates a natural language description aligned with the action. The resulting (input structure, action prompt) pairs are then fed into the LLM agent system, whose generated structure is compared against the target structure using the StructureMatcher from pymatgen to compute the desired evaluation metric.
@@
swap Swap the spatial positions of atoms at indices {index1} and {index2} in the cif file. The indices of atoms are started from 0.
@@ 3.2 AtomWorld for LLM Training
3.3 Complementary Tests
To support the analysis of AtomMotor-1K, we design a set of complementary tests spanning format literacy, spatial reasoning, and property-oriented understanding. These tests play the role of breaking down AtomMotor-1K to systematically target different levels of reasoning, and also to tease the applicability of agentic CIF modification workflows through integrating tests for perceptual skills.
@@ 3.3 Complementary Tests
      CIF-Gen: Evaluates whether the model can explicitly produce syntactically valid CIFs for simple prototype crystals (e.g., sc, fcc, bcc, perovskite), thereby examining familiarity with CIF conventions and basic materials knowledge (as opposed to the open-ended CIF generation explored by Antunes et al. (2024); Chen et al. (2024); Choudhary (2024)).
@@
      Chemical Competence Score (CCS): This test assesses a model’s latent chemical knowledge by evaluating its precision in distinguishing chemically accurate from inaccurate descriptions of crystal structures. While this test is a “perceptual skills” test, we use this to measure the effect that chemistry pretraining has on LLM performance in “motor skills” tasks. Following the methodology of Bran et al. (2025), the dataset was constructed by sampling 600 unique crystal structures from the Materials Project, with corresponding descriptions generated using Robocrystallographer (Ganose and Jain, 2019). An inaccurate dataset was then created by replacing one sentence in each original description with a sentence describing a different crystal. Because the CCS is computed from the token log-likelihoods at the model’s final layer, access to these probabilities is required; this score can be calculated only for locally-run models.
@@ 4.1 Models and Parameter Ranges
Our selection of LLMs covers frontier closed models and strong open-source baselines. We chose the Qwen-3 series to test for parameter scaling effects. We also considered science-specialised LLMs; e.g. NatureLM (Xia et al., 2025), and MatterChat (Tang et al., 2025). However, these were excluded due to either unable to produce outputs in the required CIF format, or not currently accessible via public APIs or code implementations.
For tool-augmented LLMs, we designed a preliminary framework that enables interaction with tools such as Pymatgen. Although this workflow still requires refinement and future iterations may yield stronger results, the current findings already illustrate meaningful trends.
@@ 4.3 Datasets
Figure 2: Distribution of the “before” material structures used in AtomMotor-1K (total of 250 structures) by their number of atoms. Also depicted for reference is MP’s distribution and visualisations of some structures used in AtomMotor-1K.
@@
[…]
@@ 5 Results
5.1 AtomWorld: AtomMotor-1K
The main results of AtomMotor-1K, alongside complementary tests are presented in Figure 3.a. We see some separation of the AtomWorld actions into easy (change, remove, swap, add), moderate (move, move towards, insert between) and hard difficulty (delete below, rotate around) levels based on their success rates. The moderate and hard difficulty tasks constitute greater requirements of multi-step or spatial reasoning. We also notice the mean max_dist metric increase for the more difficult tasks (minus tasks not requiring structural perturbations). The max_dist distributions for each task are shown in Appendix B.3. One interesting finding is that , the super_cell task cannot be well categorised into these difficulty tiers as the success rates range from Llama3-70B’s 4% to 98% from GPT-o3 - it’s both easy (just large-scale repetition) and difficult (requires long-context output) at the same time. The parameter scaling results in Figure 3.c and d illustrate that larger models generally achieve higher success rates and smaller displacements. However, with improvements with scale being marginal with more difficult tasks, and noting that Qwen3-32B outperformes Llama3-70B across most tasks, it suggests that architectural design and training strategies play an equally important role as parameter size. We note that the failure cases appear largely random, with outputs ranging from unmodified input structures to partially modified or slightly perturbed atom positions. A detailed mechanistic analysis would require extensive human inspection and is left for future work.
Our evaluation of our tool-augmented LLM framework on AtomWorld tasks found noticeable gains in model performance. However, the gains are somewhat limited, particularly for more complex actions. Detailed results and comprehensive analysis can be found in Appendix C.
Figure 3: a. Success rate metric across AtomMotor-1K, CIF-Repair, CIF-Gen and StructProp datasets. b. Mean max_dist metric across AtomMotor-1K and CIF-Gen datasets. c, d. Parameter scaling results on Qwen3 series. The right side are some randomly sampled structures from the tested data.
@@ 5.2 PointWorld and CIF Literacy Tests
Figure 4: Visualised results of CIF-Gen task. The left side shows the number of correctly generated CIFs for each structure type. The squares marked in red indicate cases where the single correct generation is the standard prototype. The right side shows the specific 3D crystal structures for each type, where the chemical compositions in red represent the standard prototypes.
@@ Appendix A AtomWorld Setup Details
A.1 Dataset distribution
Herein, we analyze the distributions of structures collected in our benchmark in terms of elemental composition, number of atoms, number of elements, and space groups, and compare them with those from the Materials Project. The data are shown in Figure 5, Figure 6, Figure 7, and 8, respectively. The structures in our dataset are randomly sampled from the Materials Project, with the number of atoms ranging from 10 to 100. Note that our data generator is not limited to the sampled subset; in principle, any structure file can be used for data generation, and it is capable of automatically generating an unlimited amount of benchmarking data.
Figure 5: Distribution of structures by elemental composition, compared with the Materials Project. The orange bars represent the data used in the AtomWorld Bench, while the blue bars represent those of the Materials Project.
Figure 6: Distribution of structures by the number of atoms, ranging from 10 to 100, compared with the Materials Project.
Figure 7: Distribution of the number of elements per structure compared with the Materials Project.
Figure 8: Distribution of structures by space group, compared with the Materials Project.
A.2 Data generator parameters for each action
Each action is associated with a set of parameters that specify which atoms the action targets, as well as the magnitude and manner in which the action is applied. Our data generator takes a given structure and a specified action type as input, and produces a set of randomly initialized parameters that define the action. The generator then applies the action to the structure, producing the resulting data. Table 5 lists the parameters required for each action used in this work, along with the ranges from which the data generator randomly samples their initial values.
Table 5: Parameter ranges for random actions in the data generator. The input structure contains NN atoms, and the lattice matrix is $A=(\bm{a}_{1},\bm{a}_{2},\bm{a}_{3})$.
Action Sampling ranges of parameters
change index: $[0,N)$
symbol: {H, He, …, Os}
remove index: $[0,N)$
add position: $A\bm{u},\;\bm{u}\in[0,1)^{3}$
symbol: {H, He, …, Os}
move index: $[0,N)$
displacement: $\mathcal{N}(0,\sigma^{2}I_{3}),\;\sigma=2$
move_towards index1, index2: $0,N),\;\texttt{index1}\neq\texttt{index2}$
distance: $[0.1,3)$ Å
insert_between index1, index2: $0,N),\;\texttt{index1}\neq\texttt{index2}$
symbol: {H, He, …, Os}
distance_ratio: $[0.1,0.9)$
swap index1, index2: $0,N),\;\texttt{index1}\neq\texttt{index2}$
Only between atoms with different symbols.
delete_below index: $[0,N)$
include_self: {True, False}
rotate_around index: $[0,N)$
radius: $[1.0,4.0)$ Å, capped by the structure size.
angle: $45^{\circ},315^{\circ})$
axis: $\{\pm\hat{\bm{x}},\pm\hat{\bm{y}},\pm\hat{\bm{z}}\}$
super_cell size: $(a,b,c)\in\{1,2,3,4\}^{3},a\times b\times c\leq 8,(a,b,c)\neq(1,1,1)$
A.3 Supported action prompts for PointWorld.
@@ A.2 Data generator parameters for each action
A.4 Full Prompt Templates
@@ 2. Lattice Parameter Adjustment;
A.5 Illustrative example of the framework
@@
A.6 Logic on generating CIF-Repair task
@@
A.7 DFT computation details
All density functional theory (DFT) calculations, including band gap and bulk modulus evaluations, were performed using the Vienna Ab initio Simulation Package (VASP) with the projector-augmented wave (PAW) method (Kresse and Hafner, 1993; Kresse and Furthmüller, 1996a; b; Kresse and Joubert, 1999) and the PBEsol exchange–correlation functional(Perdew et al., 2008). High-throughput workflows for both properties were automated using the atomate2 package (Ganose et al., 2025). Unless otherwise specified, calculation parameters followed the default settings in atomate2. Example calculation scripts are provided in the github repository.
@@ A.7 DFT computation details
Appendix B Supplementary Experimental Data
B.1 Prompt choice and misinterpretation
During tests, we found that certain prompt formulations could cause LLMs to misinterpret spatial actions as text-level editing, specifically for the swap action (Listing 6). Prompts that focused on textual “indices” without explicitly framing the task as a spatial transformation caused some models to attempt CIF-text rewriting instead of manipulating atomic coordinates. To verify whether this was an intrinsic ambiguity in the wording, we asked three domain experts in materials modelling to perform the same tasks using only the less explicit prompt. Both interpreted the action correctly and did not exhibit the LLM-style misinterpretation. As domain experts with long-term experience in atomic modelling, their prior exposure to structure manipulation likely provides an additional intuition for identifying the spatially intended reading, rather than a purely textual one. In contrast, LLMs appear to exhibit a default “interpretive bias” toward textual position unless the spatial nature of the task is made explicit.
Our goal, however, was not to optimize prompts, but to ensure consistent and unambiguous task interpretation. All experiments therefore use a single unified prompt set, chosen simply to remove obvious sources of misunderstanding while maintaining comparability across tasks. These prompts are likely not globally optimal but effectively prevent semantic confusion without engaging in extensive prompt tuning.
Listing 6: Less explicit and explicit spatial prompts for swap action
Less explicit prompt:
"Swap atoms at indices {self.index1} and {self.index2} in the cif file. The indices of atoms are started from 0."
Success rate (Deepseek-chat): 22%
Success rate (Qwen3 32B): 50%
Explicit spatial prompt:
"Swap the spatial positions of atoms at indices {self.index1} and {self.index2} in the cif file. The indices of atoms are started from 0."
Success rate (Deepseek-chat): 64%
Success rate (Qwen3 32B): 98%
B.2 Performance sensitivity analysis
B.2.1 Sensitivity on the number of atoms
To quantitatively investigate the influence of the number of atoms on the action success rate, we prepared test data that is as clean as possible, using a single action with consistent and simple action parameters. For the structures, we employed different supercell sizes of the same base structure. For each structure, we have tested 10 times. In this case, we chose the insert_between action to insert a Hydrogen atom in the middle between atoms at indices 3 and 5. We chose BiOF (mp-762304, P21/c, 12 atoms per unit cell) as the test system, and generated supercells with 12, 24, 48, 96, 144, and 216 atoms. This system was selected to provide moderate structural complexity while being representative of a three-element compound with ample data in the Materials Project.
Figure 10: The relation between success rate and the number of atoms (directly related to token lengths).
As shown in Figure 10, the success rates of all tested models decrease as the number of atoms in the CIF increases. The larger model (Qwen3 32B) shows a slower decline compared to the smaller ones (Qwen3 4B). Beyond 200 atoms, all three models failed in every test case. This trend suggests that while larger models are more robust to increasing input size, extremely large structures still exceed the models’ effective reasoning capacity.
B.2.2 Sensitivity on Bravais lattice types
Figure 11: The success rate under the 14 Bravais lattice types.
We also investigated the effect of structural symmetry on the action success rate. Given that there are 230 space groups in total, analyzing all of them would be impractical. Therefore, we focused on the 14 Bravais lattices for statistical analysis. To minimize the influence of other factors, we performed random sampling from the Materials Project, selecting conventional cells containing 12 to 28 atoms, and collected 10 structures for each lattice type. As shown in Figure 11, the Bravais lattices are arranged roughly in decreasing order of symmetry, from face-centered cubic (cF) to simple primitive (aP). The success rates of the three tested models show no clear correlation with the lattice symmetry. However, the models tend to achieve slightly lower accuracy on low-symmetry systems, although a definitive conclusion would require more detailed investigation.
B.2.3 Sensitivity on the action positions
We further evaluated whether the position of the target atom in the CIF sequence affects the action success rate. To this end, we prepared paired test cases using the same structure, where the target atoms appeared either early or late in the CIF atom list. All other factors were kept fixed. Specifically, we used the insert_between action with a distance_ratio of 0.45 and fixed the inserted atom type as Hydrogen. For the “early” setting, the selected atom indices were 0 and 1, whereas for the “late” setting, the last two indices in the CIF were used. The underlying structures are identical to those used in the standard insert_between tasks.
As shown in Table 7, the tested models consistently achieve higher success rates when the target atoms appear earlier in the CIF sequence. Although the difference is not large, the trend is observed across the two evaluated models. Overall, the results indicate that atom ordering has some impact on performance, but the effect remains limited under our testing conditions. To overcome this issue, we believe that the future agentic systems should use specific tools or modalities to understand the structures, then perform the actions.
Table 7: The success rate for different action positions.
Action position Deepseek-chat (%) Qwen3 32B (%)
Index 0, 1 90 96
Index -2, -1 82 80
B.2.4 Performance on moving all atoms
We further evaluated the move_all action, which requires the model to translate the entire structure by a specified displacement vector. As summarized in Table 8, the success rates of all tested models drop substantially compared to the single-atom move tasks. In many failed cases, the generated structures show noticeable randomness in the atomic positions, indicating that the models have difficulty performing consistent global translations.
Unlike the other main metrics in AtomWorld, evaluating move_all poses an additional challenge: a rigid translation preserves the structure up to translational symmetry. As a result, the standard StructureMatcher cannot be directly applied because it intentionally factors out translational degrees of freedom. To quantify deviations, we implemented a custom metric similar to StructureMatcher but without enforcing translational invariance, comparing atoms by their index correspondence and computing the resulting coordinate deviations.
Table 8: The metrics for the move_all action.
Models Success rate (%) mean max_dist (Å)
Deepseek-chat 54 0.1111
Qwen3 4B 12 0.1118
Qwen3 8B 24 0.0252
Qwen3 14B 40 0.0467
Qwen3 32B 46 0.0487
B.3 The max_dist violin plots
Appendix C Evaluations of tool augmented LLM for AtomWorld